# Effective Two-Stage Knowledge Transfer for Multi-Entity Cross-Domain Recommendation

### v1

@@ Abstract.
In recent years, the recommendation content on e-commerce platforms has become increasingly rich — a single user feed may contain multiple entities, such as selling products, short videos, and content posts. To deal with the multi-entity recommendation problem, an intuitive solution is to adopt the shared-network-based architecture for joint training. The underlying idea is to transfer the extracted knowledge from one type of entity (source entity) to another (target entity). However, different from the conventional same-entity cross-domain recommendation, multi-entity knowledge transfer encounters several important challenges: (1) data distributions of the source entity and target entity are naturally different, making the shared-network-based joint training susceptible to the negative transfer issue, (2) more importantly, the corresponding feature schema of each entity is not exactly aligned (e.g., price is an essential feature for selling product while missing for content posts), making the existing approaches no longer appropriate. Recent researchers have also experimented with the pre-training and fine-tuning paradigm. Again, they only take into account the scenarios with the same entity type and feature systems, which is inappropriate in our case. To this end, we design a pre-training & fine-tuning based Multi-entity Knowledge Transfer framework called MKT. MKT utilizes a multi-entity pre-training module to extract transferable knowledge across different entities. In particular, a feature alignment module is first applied to scale and align different feature schemas. Afterward, a couple of knowledge extractors are employed to extract the common and entity-specific knowledge. In the end, the extracted common knowledge is adopted for target entity model training. Through extensive offline and online experiments, we demonstrated the superiority of MKT over multiple State-Of-The-Art methods. MKT has also been deployed for content post recommendations on our production system.
@@ 1. Introduction
Due to the limited exposure, building an independent model (Cheng et al., 2016; Guo et al., 2017; Zhou et al., 2018, 2019; Ouyang et al., 2019) solely on the impression data of content posts can be inferior in model performance (Zhang et al., 2019). An intuitive solution is to utilize abundant user interaction information from the other entity (source domain) (Zhu et al., 2022; Khan et al., 2017), i.e. the selling products for Xianyu, and assist in the training of the content post (target domain) recommendation algorithm. Despite being promising, an effective transfer of knowledge across different entities faces several important challenges: (1) data distributions of the source entity and target entity are naturally different, making the shared-network-based joint training susceptible to the negative transfer issue, and (2) more importantly, the feature schema of different entities are not exactly aligned. For instance, price is an important feature for products while it is unavailable for content posts. This makes the conventional same-entity cross-domain recommendation approaches (Zhang et al., 2022; Liu et al., 2023) inappropriate.
@@
Indeed, most of the existing algorithms on cross-domain recommendation only handle the same entity type, in which the features of each recommendation item remain the same across different domains. Meanwhile, we do find a few studies that tackle the recommendation with multiple entities (Ouyang et al., 2020; Hu et al., 2018; Li and Tuzhilin, 2020, 2021); however, they either lack an explicit feature alignment module or fail to provide a model architecture differentiating shared and specific knowledge, both can result in the negative transfer issue. To the best of our knowledge, current research on cross-domain recommendation can be broadly divided into two groups: the joint training (Elkahky et al., 2015; Li and Tuzhilin, 2020; Hu et al., 2018; Ma et al., 2018; Tang et al., 2020; Chang et al., 2023; Zhang et al., 2023; Zhao et al., 2020) approach in which data from source and target domains is trained together within a single stage, and the pre-training & fine-tuning (Chen et al., 2021; Zhang et al., 2022; Yang et al., 2022; Liu et al., 2023; Huan et al., 2023) approach with the first stage on pre-training the source domain and the second stage on fine-tuning the target domain.
A joint training algorithm typically involves designing shared and independent parameters across domains to capture shared and unique knowledge, along with various gating architectures to dynamically regulate outputs from different domains. The goal is to transfer knowledge from the source domain to the target domain through shared parameters. Take STAR (Sheng et al., 2021) for example, Sheng et al. developed the Star Topology Adaptive Recommender (STAR) algorithm, which comprises both the shared centered parameters and domain-specific parameters. The shared parameters are designed to learn the common characteristics across domains, while the domain-specific parameters capture domain differences. However, the shared-parameter architecture often encounters the gradient conflict problem during back-propagation. This is more prominent once the training data is highly skewed towards one domain. In our production system, the amount of data for the product entity is 9 times over the content post entity, this may cause the model parameters to be governed by the product entity (Crawshaw, 2020; Xie et al., 2022; Zhang and Yang, 2021).
To resolve the negative transfer issue, researchers have experimented with the way of pre-training & fine-tuning. During pre-training, a base model is trained with an extensive amount of source data. Afterward, the obtained model or a new target model is tuned with the data from the target domain to better accommodate the target distribution. For instance, KEEP (Zhang et al., 2022) adopts a two-stage framework that comprises a supervised pre-training knowledge extraction module on a super-domain, and a plug-in network that incorporates the extracted knowledge into the target domain. However, its pre-training stage requires a vast amount of data and produces a relatively static representation of user-item interest. CTNet (Liu et al., 2023) further proposes a continuous knowledge transfer algorithm from the temporal perspective. However, it is incapable of handling entities with different feature schemas, which will be resolved in this paper.
@@
To address the problem of multi-entity knowledge transfer with heterogeneous feature schema, we propose an effective Multi-entity Knowledge Transfer algorithm (MKT for short). [^2] Note that, for simplicity, we referred to the product as the source entity and the content post as the target entity in this paper. MKT follows the pre-training & fine-tuning paradigm.
@@
During pre-training, MKT extracts useful knowledge from the source entity through training a multi-entity compatible base model. Due to the feature alignment issue, pre-training purely based on the source entity data, as done in the previous cross-domain research (Zhang et al., 2022; Liu et al., 2023), is insufficient. To better accommodate for the second-stage fine-tuning, MKT adopts the mixed data from both source and target entities for pre-training and further develops a Heterogeneous Feature Alignment (HFA) module to align the heterogeneous feature schema of the two entities. In addition, MKT employs a Common Knowledge Extractor (CKE) and an Independent Knowledge Extractor (IKE) to extract the common and independent knowledge between two entities. To better capture the semantic meaning of different knowledge representations, MKT further designs an auxiliary task to enforce the common knowledge and entity-specific knowledge to be dissimilar. The resulting multi-entity base model is then plugged into the target domain model for fine-tuning. This ultimately enhances the prediction performance for the target model (i.e., the content post recommendation model).
@@
  • By diving into the multi-entity recommendation problem, we discover several important issues and develop the corresponding solutions. Specifically, we develop a Heterogeneous Feature Alignment (HFA) module to handle different feature schemas from multiple entities and propose an auxiliary task with a polarization distribution to better extract common knowledge across entities.
  • We verify the effectiveness of MKT in industrial datasets and examine the model performance through online A/B testing on the Xianyu platform. Experimental results demonstrate that MKT outperforms the State-Of-The-Art algorithms by a fair margin, resulting in a 4.1% increase in click-through rate and a 7.1% increase in user engagement metric in online experiments.
@@ 2. Related Work
The multi-entity recommendation problem described in this paper falls under the research direction of Cross-Domain Recommendation (CDR for short); thereby, we mainly summarize the related research studies in this area. The CDR research aims to utilize the abundant data from the source domain to enhance the recommendation performance of the target domain (Fernández-Tobías et al., 2012; Zang et al., 2022; Zhu et al., 2021). This is widely recognized as an effective method to address the data sparsity and cold start issues for the target domain. Existing cross-domain algorithms can be divided into two groups: the joint training approach and the pre-training & fine-tuning approach.
@@
Considering that users are normally consistent across domains, many studies have investigated the ways of propagating user interest to different domains. For instance, MVDNN (Elkahky et al., 2015) designed a shared user tower and multiple independent item towers, that helped broadcast user knowledge across domains. MiNet (Ouyang et al., 2020) divided cross-domain user interests into long-term interest and short-term interest, and jointly models three types of user interests with an attention mechanism. However, a pure transfer of user profile knowledge is insufficient, researchers have also examined the ways of knowledge extraction from user-item feature interaction.
MMOE (Ma et al., 2018) utilized multiple shared expert networks and domain-specific gating networks to capture the commonality and difference in user-entity interaction information. PLE (Tang et al., 2020) added independent expert networks and exploited a progressive routing mechanism to extract shared and task-specific knowledge. CoNet (Hu et al., 2018) adopted cross-connection units for knowledge transfer across different domains. STAR (Sheng et al., 2021) proposed a star-shaped topology compromising globally shared parameters and domain-specific parameters, and the final domain parameters were computed with the dot product of the two parameters. PEPnet (Chang et al., 2023) designed a dynamic parameter generator that is shared across multiple domains.
@@
However, joint training algorithms inevitably face gradient conflict problems, and since the number of data samples in the source domain is usually much larger than that in the target domain, the final model is dominated by the source domain; therefore, the final model does not fully adapt to the target domain. Furthermore, joint training algorithms cannot effectively handle multi-entity problems due to the misalignment of feature patterns.
CDR with Pre-training & Fine-tuning. The pre-training & fine-tuning approach is another main solution to the cross-domain recommendation problem. Conventional methods (Chen et al., 2021; Yang et al., 2022) usually adopted the same model for both pre-training and fine-tuning. That is, we first pre-train on the source domain to obtain a base model and then fine-tune the resulting base model with the data from the target domain. However, as mentioned in previous studies (He et al., 2021; Kumar et al., 2022), those methods can easily get stuck in the local optima once the target data distribution is far away from the source data distribution.
@@
Therefore, later studies such as KEEP (Zhang et al., 2022) adopted the knowledge plugging framework, in which the pre-trained knowledge is inserted into the target domain. However, due to the large amount of training data and the corresponding model architecture, the extracted knowledge from KEEP is rather static. CTNet (Liu et al., 2023) further proposed a continuous knowledge transfer algorithm for dynamic knowledge transfer. The above knowledge plugging paradigm is indeed capable of adapting the final model to the target domain through fine-tuning; however, they can not deal with entities with different feature schemas.
@@ 3. Methodology
This section introduces the proposed MKT algorithm. We start with defining the problem and then describe the algorithmic framework in Section 3.1. In the next few sections, we will provide more details on each module of the proposed MKT algorithm. Note that the notations in this paper follow the below rules: the bold uppercase letters define matrices, the bold lowercase letters define vectors, and regular lowercase letters define scalars. All of the vectors are in the form of column vectors.
We define our research topic as the multi-entity cross-domain recommendation, in which we attempt to utilize rich user interaction data from the source entity to enhance the recommendation experience for the target entity. To clarify, in this paper, the source entity refers to the products, and the target entity refers to the content posts. Both of them are displayed to the end-users on the homepage feeds of Xianyu, while the number of impressions for products is 9 times more than the posts. In addition, if not specially mentioned, the recommendation algorithm mainly implies the Click-Through Rate (CTR) prediction model as it is the main ranking criterion for homepage recommendation.
Formally, let us denote the source entity data as $(X^{s},Y^{s})$ and the target entity data as $(X^{t},Y^{t})$, where $X$ represents the input features and $Y$ represents the corresponding label information, e.g. click. Both of them will be exploited for knowledge pre-training, and the extracted knowledge will be further applied to improve the CTR prediction performance for the content post recommendation.
@@ 3.1. Model Overview
Considering the feature difference, MEM develops a Heterogeneous Feature Alignment (HFA) module to screen and align different feature schemas, facilitating the knowledge extraction process in the subsequent Common Knowledge Extractor (CKE) module. To ensure CKE functions as intended, we introduce two independent knowledge extractors to focus on extracting entity-specific knowledge, and an auxiliary loss to enforce the entity-specific knowledge to be away from the common knowledge.
In the second stage, the low-level feature embeddings, the extracted high-level common knowledge, and the Target Entity Model are all trained together to predict user clicks on the target entity. Instead of directly concatenating the learned representations, we introduce several gating layers to better accommodate the target entity model. The resulting model will then be used for online serving. The whole process is also visualized in Figure 2.
Figure 2. An overview of the model architecture for MKT, which consists of a multi-entity base model (left) that is trained on the mixed-domain data, and a target entity model (right) for the target entity recommendation. Modules in the gray color denote the online serving part.
@@ 3.2. Multi-Entity Model Pre-training
In the below sections, we provide more details on pre-training the multi-entity model with mixed source and target entity data.
@@ 3.2.1. Feature Embedding
Here, $\mathbf{v}_{shared}^{s}$ and $\mathbf{v}_{spec}^{s}$ correspond to the embedding vectors of entity-shared features and entity-specific features respectively. $\mathbf{v}_{seq}^{s}$ is obtained through the attention mechanism. Since attention is not the key design of this paper, we omit the detail in this section. It can be Multi-Head Attention (Vaswani et al., 2017) or other sequence encoders. Features employed in the attention mechanism are part of entity-shared features, so multiple entities can share parameters in the sequence encoder. The symbol ||||| | denotes the vector concatenation operation. Similarly, for target domain sample $\mathbf{x}^{t}$, we can obtain $\mathbb{v}^{t}\in\mathbb{R}^{d_{t}\times 1}$ in the same manner.
@@ 3.2.2. Heterogeneous Feature Alignment(HFA)
In practice, we have observed that the entity-specific features significantly affect the model performance. However, simply mapping those features onto the same dimension and then sharing them across different entities is sub-optimal. With heterogeneous feature schemas, we have to design a module that can properly handle the alignment of entity-specific features. To this end, we develop a Heterogeneous Feature Alignment (HFA) structure that is capable of selecting and aligning those entity-specific features.
HFA first utilizes explicit and implicit interactions between entity-specific features and entity-shared features to jointly assess the importance of each entity-specific feature. With the importance score, HFA further augments those features with high scores. Finally, HFA maps the entity-specific features (from both source and target entities) into the same dimension for alignment, and then concatenates the resulting vector with the shared features. The whole process can be illustrated as Figure 3.
@@
For $\mathbf{v}^{s}_{spec}$, we adopt the below Formula 4 to obtain the explicit feature cross (Wang et al., 2017). Here, $\mathbf{w}_{ex,l}^{s}\in\mathbb{R}^{d_{s}\times 1}$ and $\mathbf{b}_{ex,l}^{s}\in\mathbb{R}^{d_{s}\times 1}$ represent the weight and bias for the explicit feature interaction at the $l$-th layer (note that we use ex to denote explicit). In this paper, we employ two layers of explicit feature interaction for obtaining $\mathbf{v}_{ex,2}^{s}$. The same process can be easily applied to the target entity.
(3) $\mathbb{v}_{ex,0}^{s}=\mathbb{v}^{s}_{spec}$
(4) $\mathbf{v}_{ex,l+1}^{s}=\mathbf{v}_{ex,l}^{s}{\mathbf{v}_{ex,0}^{s}}^{T}\mathbf{w}_{ex,l}^{s}+\mathbf{b}_{ex,l}^{s}+\mathbf{v}_{ex,0}^{s},l=0,1,2...$
As for the implicit feature interaction (i.e., $\mathbf{v}^{s}_{im}$ and $\mathbf{v}^{t}_{im}$), we exploit a fully connected (FC) layer and the ReLU activation function, as shown in Formula 5, for such a purpose. Here, we use im to denote the word implicit.
(5) $\mathbf{v}^{s}_{im}=ReLU(\mathbf{W}^{s}\mathbf{v}^{s}_{spec}+\mathbf{b}^{s})$
We then concatenate the resulting vectors from the explicit feature cross and implicit feature cross, and exploit an FC layer with a Sigmoid activation function to map them to vectors whose length corresponds to the total number of entity-specific features. This creates the importance score for each feature in the source entity $\mathbf{p}^{s}\in\mathbb{R}^{n_{spec}^{s}\times 1}$, and the target entity $\mathbf{p}^{t}\in\mathbb{R}^{n_{spec}^{t}\times 1}$.
(6) $\mathbb{p}^{s}=2*Sigmoid(FC(Concat(\mathbf{v}_{ex,2}^{s},\mathbb{v}^{s}_{im})))$
Based on the feature importance score, HFA then amplifies important features and down-weights the unimportant ones. In the end, the processed entity-specific features and entity-shared features are concatenated for the use of the next module. The alignment of the source and target entity features is achieved through a multilayer perceptron (MLP), which is shown below:
@@
(8) $\mathbb{q}^{s}=Concat(\mathbb{v}_{shared}^{s},ReLU(FC(Flatten(Multiply(\mathbb{V}^{s},\mathbb{p}^{s})))))$
where $Multiply$ represents element-wise multiplication, and $Flatten$ is the operation of flattening a matrix into a column vector. Finally, we obtain output vectors of the source entity and target entity with the same dimension, denoted as $\mathbb{q}^{s}\in\mathbb{R}^{d_{HFA}\times 1}$ and $\mathbb{q}^{t}\in\mathbb{R}^{d_{HFA}\times 1}$ ($d_{HFA}$ is the hyper-parameter for HFA).

### v2

@@ Abstract.
In recent years, the recommendation content on e-commerce platforms has become increasingly rich — a single user feed may contain multiple entities, such as selling products, short videos, and content posts. To deal with the multi-entity cross-domain recommendation problem, an intuitive solution is to adopt the shared-network-based architecture for joint training. The underlying idea is to transfer the knowledge from one type of entity (source entity) to another (target entity). However, different from the conventional same-entity cross-domain recommendation, multi-entity knowledge transfer encounters several important challenges: (1) data distributions of the source entity and target entity are naturally different, making the shared-network-based joint training susceptible to the negative transfer issue, (2) the corresponding feature schema of each entity is not exactly aligned (e.g., price is an essential feature for selling product while missing for content posts), making the existing approaches no longer appropriate. Recent researchers have also experimented with the pre-training and fine-tuning paradigm. Again, they only take into account the scenarios with the same entity type and feature schemas, which is inappropriate in our case. To this end, we design a pre-training & fine-tuning based Multi-entity Knowledge Transfer framework called MKT. MKT utilizes a multi-entity pre-training module to extract transferable knowledge across different entities. In particular, a feature alignment module is first applied to scale and align different feature schemas. Afterward, a couple of knowledge extractors are employed to extract the common and independent knowledge. In the end, the extracted common knowledge is adopted for target entity model training. Through extensive offline and online experiments on public and industrial datasets, we demonstrated the superiority of MKT over multiple State-Of-The-Art methods. MKT has also been deployed for content post recommendations on our production system.
@@ 1. Introduction
Due to the limited exposure, building an independent model (Cheng et al., 2016; Guo et al., 2017; Zhou et al., 2018, 2019; Ouyang et al., 2019) solely on the scarce impression data of content posts can be inferior in model performance (Zhang et al., 2019). An intuitive solution is to utilize abundant user interaction information from the other entity (source domain) (Zhu et al., 2022; Khan et al., 2017), i.e. the selling products for Xianyu, and assist in the training of the content post (target domain) recommendation algorithm. Despite being promising, an effective transfer of knowledge across different entities faces several important challenges: (1) data distributions of the source entity and target entity are naturally different, making the shared-network-based joint training susceptible to the negative transfer issue, and (2) the feature schemas of different entities are not exactly aligned. For instance, price is an important feature for products while it is unavailable for content posts. This makes the conventional cross-domain recommendation approaches (Zhang et al., 2022; Liu et al., 2023) inappropriate.
@@
To the best of our knowledge, current research on cross-domain recommendation can be broadly divided into two groups: the joint training (Elkahky et al., 2015; Li and Tuzhilin, 2020; Hu et al., 2018; Ma et al., 2018; Tang et al., 2020; Chang et al., 2023; Zhang et al., 2023; Zhao et al., 2020; Ouyang et al., 2020; Li and Tuzhilin, 2021) approach in which data from source and target domains is trained together within a single stage, and the pre-training & fine-tuning (Chen et al., 2021; Zhang et al., 2022; Yang et al., 2022; Liu et al., 2023; Huan et al., 2023) approach with the first stage on pre-training the source domain and the second stage on fine-tuning the target domain. Joint training algorithms typically transfer knowledge from the source domain to the target domain by designing shared parameters, most of them only consider the same entity type (Ma et al., 2018; Chang et al., 2023; Sheng et al., 2021), while a few studies consider multiple entities (Ouyang et al., 2020; Hu et al., 2018; Li and Tuzhilin, 2020). However, the shared-parameter architecture often encounters the negative transfer (Zhang and Yang, 2021) problem during back-propagation. This is more prominent once the training data is highly skewed towards one domain. In our production system, the amount of data for the product entity is 3 times over the content post entity, this may cause the model parameters to be governed by the product entity (Crawshaw, 2020; Xie et al., 2022). To address negative transfer problem, researchers have tried pre-training and fine-tuning. First, a base model is trained using a lot of source data. Then, the obtained model is adjusted with target domain data, or a new target model is tuned with the knowledge from the obtained model to fit the target distribution better (Zhang et al., 2022; Liu et al., 2023). However, they only take into account the scenarios with the same entity type and feature schema.
@@
To address the problem of multi-entity knowledge transfer with heterogeneous feature schema, we propose an effective Multi-entity Knowledge Transfer algorithm (MKT for short). MKT follows the pre-training & fine-tuning paradigm.
@@
During pre-training, MKT extracts useful knowledge from the source entity through training a multi-entity compatible base model. To better accommodate for the second-stage fine-tuning, MKT adopts the mixed data from both source and target entities for pre-training and further develops a Heterogeneous Feature Alignment (HFA) module to align the heterogeneous feature schema of the two entities. In addition, MKT employs one Common Knowledge Extractor (CKE) and two Independent Knowledge Extractor (IKE) to extract the common and independent knowledge between two entities. To better capture the semantic meaning of different knowledge representations, MKT further designs an auxiliary task to enforce the common knowledge and independent knowledge to be dissimilar. The resulting multi-entity base model is then plugged into the target domain model for fine-tuning. MKT performs hierarchical transfer of knowledge with different semantics, including user features, entity-shared features, sequence features and user-entity cross vector, and uses gate to filter the knowledge from the pre-training stage to obtain more suitable parts. This ultimately enhances the prediction performance for the target model (i.e., the content post recommendation model).
@@
  • By diving into the multi-entity recommendation problem, we discover several important issues and develop the corresponding solutions. Specifically, we develop a Heterogeneous Feature Alignment (HFA) module to handle different feature schemas from multiple entities and propose an auxiliary task to better extract common knowledge across entities. We perform hierarchical transfer of knowledge with different semantics and use gate to filter them.
  • We verify the effectiveness of MKT on public and industrial datasets and examine the model performance through online A/B testing on the Xianyu platform. Experimental results demonstrate that MKT outperforms the State-Of-The-Art algorithms by a fair margin, resulting in a 4.1% increase in click-through rate and a 7.1% increase in user engagement metric in online experiments.
@@ 2. Related Work
The multi-entity recommendation problem described in this paper falls under the research direction of Cross-Domain Recommendation (CDR for short). The CDR research aims to utilize the abundant data from the source domain to enhance the recommendation performance of the target domain (Fernández-Tobías et al., 2012; Zang et al., 2022; Zhu et al., 2021). Existing cross-domain algorithms can be divided into two groups: the joint training approach and the pre-training & fine-tuning approach.
@@
Considering that users are normally consistent across domains, many studies have investigated the ways of propagating user interest to different domains. For instance, MVDNN (Elkahky et al., 2015) designed a shared user tower and multiple independent item towers, that helped broadcast user knowledge across domains. MiNet (Ouyang et al., 2020) divided cross-domain user interests into long-term interest and short-term interest, and jointly models three types of user interests with an attention mechanism. However, a pure transfer of user profile knowledge is insufficient, researchers have also examined the ways of knowledge extraction from user-item feature interaction. MMOE (Ma et al., 2018) utilized multiple shared expert networks and domain-specific gating networks to capture the commonality and difference in user-entity interaction information. PLE (Tang et al., 2020) added independent expert networks and exploited a progressive routing mechanism to extract shared and task-specific knowledge. CoNet (Hu et al., 2018) adopted cross-connection units for knowledge transfer across different domains. STAR (Sheng et al., 2021) proposed a star-shaped topology compromising globally shared parameters and domain-specific parameters, and the final domain parameters were computed with the dot product of the two parameters. PEPNet (Chang et al., 2023) designed a dynamic parameter generator that is shared across multiple domains.
@@
However, joint training algorithms inevitably face gradient conflict problems, and since the number of data samples in the source domain is usually much larger than that in the target domain, the final model is dominated by the source domain; therefore, the final model does not fully adapt to the target domain.
@@
CDR with Pre-training & Fine-tuning. The pre-training & fine-tuning approach is another main solution to the cross-domain recommendation problem. Conventional methods (Chen et al., 2021; Yang et al., 2022) usually adopted the same model for both pre-training and fine-tuning. That is, we first pre-train on the source domain to obtain a base model and then fine-tune the resulting base model with the data from the target domain. However, as mentioned in previous studies (He et al., 2021; Kumar et al., 2022), those methods can easily get stuck in the local optima once the target data distribution is far away from the source data distribution. Therefore, later studies such as KEEP (Zhang et al., 2022) adopted the knowledge plugging framework, in which the pre-trained knowledge is inserted into the target domain. However, due to the large amount of training data and the corresponding model architecture, the extracted knowledge from KEEP is rather static. CTNet (Liu et al., 2023) further proposed a continuous knowledge transfer algorithm for dynamic knowledge transfer. The above knowledge plugging paradigm is indeed capable of adapting the final model to the target domain through fine-tuning; however, they can not deal with entities with different feature schemas.
@@ 3. Methodology
This section introduces the proposed MKT algorithm. We define our research topic as the multi-entity cross-domain recommendation, in which we attempt to utilize rich user interaction data from the source entity to enhance the Click-Through Rate (CTR) for the target entity. Formally, let us denote the source entity data as $(X^{s},Y^{s})$ and the target entity data as $(X^{t},Y^{t})$, where $X$ represents the input features and $Y$ represents the click label information. Note that the notations in this paper follow the below rules: the bold uppercase letters define matrices, the bold lowercase letters define vectors, and regular lowercase letters define scalars. All of the vectors are in the form of column vectors.
@@ 3.1. Model Overview
Considering the feature difference, MEM develops a Heterogeneous Feature Alignment (HFA) module to scale and align different feature schemas, facilitating the knowledge extraction process in the subsequent Common Knowledge Extractor (CKE) module. To ensure CKE functions as intended, we introduce two independent knowledge extractors to focus on extracting the independent knowledge, and an auxiliary loss to enforce the independent knowledge to be away from the common knowledge.
In the second stage, the TEM is trained with the frozen low-level feature embeddings and extracted common knowledge of MEM to predict user clicks on the target entity. Instead of directly concatenating the learned representations, we introduce the gate layers to better accommodate the target entity model. The whole process is also visualized in Figure 2.
Figure 2. An overview of the model architecture for MKT, which consists of a multi-entity model (left) that is trained on the mixed-domain data, and a target entity model (right) for the target entity recommendation. Modules in the gray color denote the online serving part.
@@ 3.2. Multi-Entity Model Pre-training
In the below sections, we provide more details on pre-training the MEM with mixed source and target entity data.
@@ 3.2.1. Feature Embedding
Here, $\mathbf{v}_{u}^{s}$ represents the embedding vector of user features, $\mathbf{e}$ denotes the embedding vector of one feature, $n_{u}$ is the number of user features, $\mathbf{v}_{shared}^{s}$ and $\mathbf{v}_{spec}^{s}$ corresponds to the embedding vector of entity-shared features and entity-specific features respectively. $\mathbf{v}_{seq}^{s}$ is obtained through the attention mechanism. Since attention is not the key design of this paper, we omit the detail in this section. It can be Multi-Head Attention (Vaswani et al., 2017) or other sequence encoders. Features employed in the attention mechanism are part of entity-shared features, so multiple entities can share parameters in the sequence encoder. The symbol ||||| | denotes the vector concatenation operation. Similarly, for target domain sample $\mathbf{x}^{t}$, we can obtain $\mathbb{v}^{t}\in\mathbb{R}^{d_{t}\times 1}$ in the same manner.
@@ 3.2.2. Heterogeneous Feature Alignment(HFA)
In practice, entity-specific features significantly affect the model performance, some features are beneficial to the extraction of common knowledge, while others are harmful. And, with heterogeneous feature schemas, we have to design a module that can properly handle the alignment of entity-specific features. To this end, we develop a Heterogeneous Feature Alignment (HFA) structure that is capable of scaling and aligning those entity-specific features.
@@
HFA first utilizes explicit and implicit feature cross between entity-specific features and shared features (user profile, click sequence, entity-shared feature) to jointly assess the importance of each entity-specific feature. Then, HFA augments those features with the importance score. Finally, HFA maps the entity-specific features (from both source and target entities) into the same dimension for alignment, and then concatenates the resulting vector with the shared features. The whole process can be illustrated as Figure 3. Taking source domain samples as an example. For $\mathbf{v}^{s}_{spec}$, we adopt the Cross network in DCN (Wang et al., 2017) as the explicit feature cross and exploit multilayer perceptron(MLP) as the implicit feature cross. Here, we use ex and im to denote the word explicit and implicit represently.
(3) $\mathbf{v}^{s}_{ex}=DCN(\mathbf{v}^{s}_{spec}),\mathbf{v}^{s}_{im}=MLP(\mathbf{v}^{s}_{spec})$
We then concatenate the resulting vectors from the explicit feature cross and implicit feature cross, and exploit an FC layer with a Sigmoid activation function to map them to vectors whose length corresponds to the total number of entity-specific features. This creates the importance score for each feature in the source entity $\mathbf{p}^{s}\in\mathbb{R}^{n_{spec}^{s}\times 1}$.
(4) $\mathbb{p}^{s}=2*Sigmoid(FC(Concat(\mathbf{v}_{ex}^{s},\mathbb{v}^{s}_{im})))$
Based on the feature importance score, HFA then amplifies important features and down-weights the unimportant ones. In the end, the processed entity-specific features and entity-shared features are concatenated for the use of the next module. The alignment of the source and target entity features is achieved through MLP, which is shown below:
@@
(6) $\mathbb{q}^{s}=Concat(\mathbb{v}_{shared}^{s},MLP(Flatten(Multiply(\mathbb{V}^{s},\mathbb{p}^{s}))))$
where $d_{e}$ is the dimension of embedding vector of one feature, $Multiply$ represents element-wise multiplication, and $Flatten$ is the operation of flattening a matrix into a column vector. Finally, we obtain output vectors of the source entity and target entity with the same dimension, denoted as $\mathbb{q}^{s}\in\mathbb{R}^{d_{HFA}\times 1}$ and $\mathbb{q}^{t}\in\mathbb{R}^{d_{HFA}\times 1}$ ($d_{HFA}$ is the hyper-parameter for HFA).